SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Mar.21, 2006)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker

Operations

  • The downtime for the Q2 chamber and bellows installation is scheduled to take place from March 22 (Wednesday) to April 4 (Tuesday).
  • Selina will be giving a status report for TRG after the detector operation meeting at IR2 tomorrow. A link to the talk is provided here.

EMT

  • Dan will update the EMT online lib using the up-to-date release soon.

DCT/DCZ

  • Marc Kelly will be arriving at SLAC tomorrow afternoon and will be working on the TSF firmware issue during his stay.

IFT

L3

  • It looked like we had a L3 node trim in swing shift of March 16. This happened in run 63061. Rainer tried to look into the logfile to see what caused the trim, but no hint.

Data Quality and Monitoring

  • The IFR problem earlier this month triggered a discussion of marking runs flawed in the DQG meeting. Rahmat asked which group is responsible for luminosity check, but it seems we don't know who that is.
  • We need to properly label/update the GLT object names for the fastmon plots.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.